Replace with user/time stamp
Projects/ Work Flow
- total Syk protocol now in place -- problems?
- supernatants being accumulated -- long term storage?
- can we send off site?
- possible space on the 14th floor?
- re-prioritize use of existing space
- fixed cells for Treg -- Kenny working through; will need to plan for analysis
- all samples sent from other labs have been acquired
- ICATA/ BioCSI/ URECA
- emphasis on BioCSI analysis
- Have not had confirmation from IRB of GCO reciept
- Have printed new slides and will finish 'validation' this morning
- Neisha will redirect to Arah2 collaboration
- Lara out any day -- are we ready?
- set up CD25 depletion template
- contamination issues?
- ITN -- status? involvement from Julie?
Recruitment/ Team Update
- Performance reviews over next week
- review procedure for HTS installation
- discussion of centralized flow staff
- pipette calibration
Previous ToDo Items
- create protocol for total Syk measurement (wayne) -- done
- Re-organize data between the three Macs and add more back ups
New ToDo Items
- Madhan and Wayne f/u with HS and AG re: CoFAR storage
- analyze CoFAR with tSyk
- Priya will reconcile missing CoFAR data
- Alyssa will f/u with Lara re: BM1
- Wayne to discuss with JHU re: pilot samples from Mary Jane